OpenStudy (faiqraees):

Prove by mathematical induction for all positive integral values of x, x(x^2+5) is divisible by 6.

OpenStudy (freckles):

the first positive integer is 1 so have you proved it for x=1 yet?

OpenStudy (faiqraees):

@ganeshie8 @ParthKohli

OpenStudy (faiqraees):

No I actually am not aware of mathematical induction techniques.

OpenStudy (freckles):

You prove it for the base case. Then you assume it is true for some k within the chosen domain of the problem. Then you show it is true for the case x=k+1.

OpenStudy (freckles):

the base case being x=1 in this case since it is the first positive integer

OpenStudy (faiqraees):

Can you explain it in a little more detail

ganeshie8 (ganeshie8):

I know you want it proven by induction, but here is a quick way to convicne yourself that the given statement is indeed true : \[x(x^2+5)\equiv x(x^2-1) = (x-1)x(x+1)\] Recall that the product of any 3 consecutive integers is divisible by 3 factorial

OpenStudy (faiqraees):

How is +5 equals to -1

ganeshie8 (ganeshie8):

5 = 6 - 1 since 6 is dividible by 6, you can replace it by 0

ganeshie8 (ganeshie8):

Perhaps this is more clear : \[x(x^2+5) = x(x^2-1+6) =x(x^2-1) + 6x = (x-1)x(x+1) + 6x \] Again, this has nothing to do with induction...

OpenStudy (faiqraees):

@ganeshie8 Okay now can you explain it woth induction

ganeshie8 (ganeshie8):

You need to know how/why induction works first

OpenStudy (faiqraees):

Yeah

ganeshie8 (ganeshie8):

Let me ask you a question

ganeshie8 (ganeshie8):

Consider the set of negative integers. Is there a minimum element in this set ? If a minimum element exists, what is it ?

OpenStudy (faiqraees):

Yes and I suppose 1, and what's the name of the technique you used above to prove the statement

ganeshie8 (ganeshie8):

do you mean, 1 is the minimum element in the set of "positive integers"

OpenStudy (faiqraees):

Oh they are saying that you have to use positive integral values

OpenStudy (faiqraees):

and what's the name of the technique you used above to prove the statement

ganeshie8 (ganeshie8):

Earlier proof is based on simple divisibility arguments We expressed the given expression as a product of 3 consecutive integers; then we used the fact that product of any 3 consecutive integers is divisible by 3 factorial.

ganeshie8 (ganeshie8):

don't pay too much attention to it you want to do this using induction right ?

OpenStudy (faiqraees):

Yes

ganeshie8 (ganeshie8):

Lets get back to induction

OpenStudy (faiqraees):

sure

ganeshie8 (ganeshie8):

Suppose I give you a machine that does this : if you input a positive integer \(k\), the machine spits out the next integer \(k+1\)